# Role of interaction anisotropy in polymer cononsolvency: insights from the Flory-Huggins-Potts framework

### v2

@@ 2.5 Conformational analysis
For consistent comparisons across conditions, we report a normalized radius of gyration, $\frac{\langle R_{\text{g}}\rangle}{\langle R_{\text{g}}\rangle_{\text{ath}}}$, where $\langle R_{\text{g}}\rangle$ is the ensemble-averaged $R_{\text{g}}$ for a given system and $\langle{R}_{\text{g}}\rangle_{\text{ath}}$ is a reference given by an athermal simulation incorporating only excluded-volume interactions. For chains with $N_{\text{m}}=32$, we empirically find that $\langle{R}_{\text{g}}\rangle_{\text{ath}}=3.83\pm 0.003$. For simplicity, we use $\langle R_{\text{g}}\rangle_{\text{ath}}=3.83$ in out calculations.
@@ 2.6 Solvation analysis
To distinguish cononsolvency mechanisms at the single-chain level, we compare the local solvation environment of the polymer from Monte Carlo simulations with ideal-mixing predictions. Deviations are quantified by the normalized excess number of interactions
$N_{\kappa}^{\text{ex}}$ $=N_{\kappa}-N_{\kappa}^{\text{id}}$
$\frac{N_{\kappa}^{\text{ex}}}{N_{\kappa}^{\text{id}}}$ $=\frac{N_{\kappa}}{N_{\kappa}^{\text{id}}}-1=\tilde{N}_{\kappa}-1$ (13)
where $N_{\kappa}$ specifies the number of interactions under some restriction $\kappa$, and the superscripts ‘ex’ and ‘id’ indicate excess and ideal values, respectively. The restriction of $\kappa$ for our analysis relates to some pairing of specific species or particles and/or whether interactions are aligned versus misaligned.
The number of pairwise interactions between monomer, solvent, and cosolvent particles is computed as
$N_{\text{mm}}$ $=\frac{1}{2}\sum_{i\in\mathcal{P}}\sum_{j\in\mathcal{N}(i)}\mathbf{1}_{j}^{\mathcal{P}},$ (14)
$N_{\text{ms}}$ $=\sum_{i\in\mathcal{P}}\sum_{j\in\mathcal{N}(i)}\mathbf{1}_{j}^{\mathcal{S}},$ (15)
$N_{\text{mc}}$ $=\sum_{i\in\mathcal{P}}\sum_{j\in\mathcal{N}(i)}\mathbf{1}_{j}^{\mathcal{C}},$ (16)
$N_{\text{sc}}$ $=\sum_{i\in\mathcal{S}}\sum_{j\in\mathcal{N}(i)}\mathbf{1}_{j}^{\mathcal{C}}.$ (17)
where $N_{\alpha_{i}\alpha_{j}}$ indicates the number of unique interactions between species $\alpha_{i}$ and $\alpha_{j}$ and $\mathcal{P}$, $\mathcal{S}$, and $\mathcal{C}$ respectively denote the sets of all polymer, solvent, and cosolvent particles, and $\mathbf{1}_{j}^{\mathcal{X}}$ is an indicator function equal to unity if the indexed particle $j$ is in the set $\mathcal{X}$ and zero otherwise.
To count only interactions directly involved with solvation, we define the solvation shell of the polymer as the union of all non-polymer particles in its neighborhood:
$\mathcal{P}_{\text{sc}}=\bigcup_{i\in\mathcal{P}}(\mathcal{N}(i)\backslash\mathcal{P})=\mathcal{P}_{\text{s}}\cup\mathcal{P}_{\text{c}},$ (18)
where $\mathcal{P}_{\text{s}}$ and $\mathcal{P}_{\text{c}}$ now denote sets of solvent and cosolvent particles within $\mathcal{P}_{\text{sc}}$, which is the set of all solvent and cosolvent particles within the neighborhood of $\mathcal{P}$.
@@
where the last expression for the number of either monomer-solvent or monomer-cosolvent interactions exploits the restriction that the overall number of interactions involving the polymer is conserved on the lattice. For chains with $N_{\text{m}}=32$, we estimate $N_{\text{mm}}^{\text{id}}$ by computing the average number of monomer-monomer interactions from athermal simulations of a chain in solvent. These simulations yield $\langle N_{\text{mm}}\rangle_{\text{ath}}=62.91\pm 0.04$. We note that a chain with $N_{\text{m}}$ monomers must have at least $N_{\text{m}}-1$ interactions, accounting for bonded neighbors, making the result above approximately and empirically $2N_{\text{m}}-1$. For simplicity, we use $N_{\text{mm}}^{\text{id}}=63$ in all calculations. Similarly, the expected number of solvent-cosolvent interactions is approximated as
@@
We characterize the solvation shell by assigning every cosolvent in $\mathcal{P}_{c}$ to one of three classes—bridging, mediating, or frustrated—and then calculating the mean energy per particle for each class. In the anisotropic model a bridging cosolvent $\Phi_{\text{b}}^{\curlywedge}$ is oriented parallel to, and in contact with, two monomer units that lie at least three bonds apart, thereby forming an enthalpic bridge between distant chain segments (Fig. 1a). A mediating cosolvent $\Phi_{\text{m}}^{\curlywedge}$ aligns with at least one monomer and one solvent particle but does not meet the bridging criterion. All other cosolvents are labeled passive $\Phi_{\text{p}}^{\curlywedge}$ (Fig. 1b).
We also make analogous classifications for isotropic systems but remove any conditions associated with orientation, which plays no role. A bridging cosolvent $\Phi_{\text{b}}^{\circ}$ contacts two monomers separated by at least three bonds; a mediating cosolvent $\Phi_{\text{m}}^{\circ}$ contacts one monomer and one solvent particle without bridging; the remainder are passive $\Phi_{\text{p}}^{\circ}$. After classifying the various cosolvents, we report the normalized mean energy per particle for bridging and mediating cosolvents, denoted $\bar{E}_{b}$ and $\bar{E}_{m}$, respectively; the passive set provides a baseline for comparison.
@@ 3 Results and discussion
Our analysis proceeds in two parts. In Section 3.1, we use mean-field theory to identify the energetic regimes that drive cononsolvency through macroscopic phase separation. These results establish conceptual scaffolding and terminology consistent with prior literature. While similar results have appeared variously,45, 49, 60, 50, 63, 51we include this analysis for completeness since application of FHP is new, and specific results here directly inform our subsequent system selection. Readers familiar with cononsolvency may skim this section. In Section 3.2, we employ lattice Monte Carlo simulations to examine microscopic behavior. We strategically select two pairs of systems with equivalent $\chi^{\text{eff}}_{ij}$ from phase-separating regimes, but for a given pair, system differ in their Hamiltonians; one features only isotropic interactions and the other incorporates orientation-dependent terms. Through detailed analysis of coil-globule-coil transitions, which are observed for all systems, we elucidate how interaction anisotropy influences cononsolvency at the microscopic level.
Fig.  2: Classification of phase behavior for ternary polymer solutions described by the FHP framework. (a) Regions in interaction parameter space $\boldsymbol{X}^{\text{FHP}}=(\chi^{\text{FHP}}_{\text{ps}},\chi^{\text{FHP}}_{\text{pc}},\chi^{\text{FHP}}_{\text{sc}})$ where phase separation is required at some composition $\boldsymbol{\phi}$, as determined by the mean-field stability criterion ($|\mathbf{H}|<0$). Markers indicate parameter sets that yield instability; cross-sectional planes at $\chi^{\text{FHP}}_{\text{sc}}=0$, 4-4- 4, 7-7- 7, and 10-10- 10 correspond to panels (b-e). (b-e) Two-dimensional slices of (a) showing phase behavior as a function of $\chi^{\text{FHP}}_{\text{ps}}$ and $\chi^{\text{FHP}}_{\text{pc}}$ for fixed values of $\chi^{\text{FHP}}_{\text{sc}}$. Shaded regions indicate parameter combinations that guarantee phase separation at some composition. Regions of observed two-phase coexistence are labeled $\mathcal{R}_{\text{sc}}$, $\mathcal{R}_{\text{ps}}$, and $\mathcal{R}_{\text{pc}}$, according to the most negative interaction parameter. Square and diamond markers in (b) and (e) denote parameter sets selected for further analysis in Section 3.2.
@@ 3.1 Analysis of phase behavior
Consistent with prior literature using FH theory,45, 49, 60, 50, 63 there are three distinct parameter regimes characterized by the strongest interaction type present (Fig. 2). We denote these regimes as $\mathcal{R}_{ij}$, corresponding to the region of parameter space where $(\chi^{\text{FHP}}_{ij}<\chi^{\text{FHP}}_{ik},\chi^{\text{FHP}}_{jk}\leq 0)$ (i.e., the strongest attractive interaction is between species $i$ and $j$). The emergence of cononsolvency, in terms of competing interactions, can be understood by examining successive cross sections in Fig. 2a. Initially, at $\chi^{\text{FHP}}_{\text{sc}}=0$, two distinct regions, $\mathcal{R}_{\text{ps}}$ and $\mathcal{R}_{\text{pc}}$, exhibit phase separation for a broad range of parameters, for which polymer-solvent or polymer-cosolvent interactions are strongly favorable relative to interactions between solvent and cosolvent (Fig. 2b). As solvent-cosolvent interactions become increasingly favorable (more negative $\chi^{\text{FHP}}_{\text{sc}}$), the areas of $\mathcal{R}_{\text{ps}}$ and $\mathcal{R}_{\text{pc}}$ shrink, as interactions are more balanced. Concurrently, a third region, $\mathcal{R}_{\text{sc}}$, emerges where solvent-cosolvent interactions dominate. (Fig. 2c) Eventually, as solvent-cosolvent affinity strengthens further, $\mathcal{R}_{\text{ps}}$ and $\mathcal{R}_{\text{pc}}$ vanish entirely, leaving an expanded $\mathcal{R}_{\text{sc}}$ region (Fig. 2d,e). This analysis provides a baseline expectation that cononsolvency can be driven by different dominant interactions, be it polymer-solvent, polymer-cosolvent, or solvent-cosolvent, irrespective of microscopic details.
@@
Prior work has suggested that the shape, location, and extent of coexistence regions, as well as the orientation of tie lines, may reflect distinct underlying mechanisms of cononsolvency. 50 To investigate this, we constructed full ternary phase diagrams for representative parameter sets drawn from each of the three regimes: $\mathcal{R}_{\text{ps}}$, $\mathcal{R}_{\text{pc}}$, and $\mathcal{R}_{\text{sc}}$. Indeed, the shapes of binodal curves and the orientations of tie lines differ based on the dominant interaction present in the system (Supplementary Information, Fig. S1) When interactions between a polymer and one of the solvents dominate, tie lines connect phases in which one is rich in polymer and the better solvent, while the other is rich in the lesser solvent and depleted in polymer. When solvent-cosolvent interactions dominate, the tie lines orient nearly perpendicular to the polymer composition axis, leading to coexistence between a very polymer-lean phase with mixed solvent and a polymer-rich phase. Nevertheless, while these observations may serve as useful indicators of the dominant underlying interactions, they remain qualitative and do not offer insight into the microscopic details of the interactions that underlie cononsolvency.
@@ 3.2 Analysis of simulations
To address this, we perform Monte Carlo (MC) simulations to study the conformational behavior and solvation-shell characteristics of a single polymer chain in mixed solvents for four systems. In Section 3.2.1, we describe the selection of these systems and broadly categorize their behavior. The results supporting this categorization are then detailed in Sections 3.2.2 and 3.2.3.
@@ 3.2.1 Overview of systems and mechanisms
Table 1: Parameters used in Monte Carlo simulations of FHP models. Parameters for label type $\mathcal{R}_{ij}^{\circ}$ are for isotropic simulations $(\Delta_{ij}=0)$ and those for type $\mathcal{R}_{ij}^{\curlywedge}$ are for anisotropic simulations $(\Delta_{ij}\neq 0)$.
@@
$\mathcal{R}_{\text{sc}}^{\circ}$1 -2 -10 0 0
$\mathcal{R}_{\text{sc}}^{\curlywedge}$1 -2 -10 -0.5 -0.5
$\mathcal{R}_{\text{pc}}^{\circ}$1 -10 0 0 0
$\mathcal{R}_{\text{pc}}^{\curlywedge}$1 -10 0 -0.5 -0.5
In all systems, we find evidence of single-chain collapse at certain mixture compositions based on observed $R_{\text{g}}$ relative to that in either pure solvent. Based on further analysis of the solvation environment around the polymer, we then associate these chain collapses with specific mechanisms within the vernacular of the cononsolvency literature. Simulations with parameter sets that are derived from $\mathcal{R}_{\text{sc}}$ show a collapse through preferential mixing of solvent and cosolvent; this mechanism tends to yield a “dry” globule. In this case, including anisotropic interactions with cosolvent species qualitatively alters the cosolvent-induced coil-globule transition. Parameter sets from $\mathcal{R}_{\text{pc}}$ collapse through preferential adsorption of cosolvent; this mechanism tends to yield a “wet” globule. For this situation, the role of anisotropy is apparent only in nuanced changes to the polymer solvation environment in its collapsed state.
Fig.  3: A comparison of conformational characteristics and environment of a polymer from $\mathcal{R}_{\text{sc}}^{\circ}$ (blue circles) and $\mathcal{R}_{\text{sc}}^{\curlywedge}$ (orange triangles). (a) Normalized single-chain $R_{\text{g}}$ and normalized excess (b) monomer-monomer, (c) monomer-cosolvent and (d) solvent-cosolvent interactions as cosolvent fraction $x_{c}$ is varied from 0 to 111. In (a), the dashed horizontal line is a guide for the reference to a maximally compact polymer and the white arrow shows the shift in the minima. In (a) and (b), the white arrow shows a marked shift in the behavior of the curves. Horizontal axis labels are share betwen panels (a) and (c) as well as (b) and (d). Error bars correspond to the standard error of the mean and are generally smaller than the symbol size.

### v3

@@ 2.5 Conformational analysis
For consistent comparisons across conditions, we report a normalized radius of gyration, $\frac{\langle R_{\text{g}}\rangle}{\langle R_{\text{g}}\rangle_{\text{ath}}}$, where $\langle R_{\text{g}}\rangle$ is the ensemble-averaged $R_{\text{g}}$ for a given system and $\langle{R}_{\text{g}}\rangle_{\text{ath}}$ is a reference given by an athermal simulation incorporating only excluded-volume interactions. For chains with $N_{\text{m}}=72$, we empirically find that $\langle{R}_{\text{g}}\rangle_{\text{ath}}=6.15\pm 0.005$. For simplicity, we use $\langle R_{\text{g}}\rangle_{\text{ath}}=6.15$ in out calculations.
@@ 2.6 Solvation analysis
To distinguish cononsolvency mechanisms at the single‑chain level, we compare the local solvation environment of the polymer from Monte Carlo simulations with ideal‑mixing predictions. Deviations are quantified by the normalized excess number of interactions
$N_{\kappa}^{\text{ex}}$ $=N_{\kappa}-N_{\kappa}^{\text{id}}$ (14)
$\frac{N_{\kappa}^{\text{ex}}}{N_{\kappa}^{\text{id}}}$ $=\tilde{N}_{\kappa}-1$ (15)
where $N_{\kappa}$ specifies the number of interactions under some restriction $\kappa$, and the superscripts ‘ex’ and ‘id’ indicate excess and ideal values, respectively. The restriction of $\kappa$ for our analysis relates to some pairing of specific species or particles and/or whether interactions are aligned versus misaligned. In the following, numbers of unique pairwise interactions between monomer, solvent, and cosolvent particles are represented with $N_{\alpha_{i}\alpha_{j}}$ where $\alpha_{i}$ and $\alpha_{j}$ are the chemical species p, s, and c.
@@
where the last expression for the number of either monomer-solvent or monomer-cosolvent interactions exploits the restriction that the overall number of interactions involving the polymer is conserved on the lattice. For chains with $N_{\text{m}}=72$, we estimate $N_{\text{mm}}^{\text{id}}$ by computing the average number of monomer-monomer interactions from athermal simulations of a chain in solvent. These simulations yield $\langle N_{\text{mm}}\rangle_{\text{ath}}=150.12\pm 0.05$. For simplicity, we use $N_{\text{mm}}^{\text{id}}=150$ in all calculations. Similarly, the expected number of solvent-cosolvent interactions is approximated as
@@
To assess the structure of the solvation shell, we compute the fraction of ‘bridging’ cosolvents $\Phi_{\text{b}}^{\curlywedge}$, ‘mediating’ cosolvent $\Phi_{\text{m}}^{\curlywedge}$, and ‘passive’ cosolvents $\Phi_{p}^{\curlywedge}$ in the solvation shell. The solvation shell is defined as the set of solvent and cosolvent particles in direct contact with the polymer i.e., nearest, next-nearest, or next-next-neareet neighbor). The classification of any given solvent as bridging, mediating, or passive depends on its underlying Hamiltonian. For systems with orientation-dependent interactions, a bridging cosolvent is is aligned and in contact with, two monomer units that lie at least three bonds apart, thereby forming an enthalpic bridge between distant chain segments (Fig. 1a). A mediating cosolvent aligns with at least one monomer and one solvent particle but does not meet the bridging criterion. All other cosolvents are labeled passive (Fig. 1b). For systems with only isotropic interactions, we make analogous classifications based on the proximity and composition of the solvation shell, but the conditions regarding orienation are relaxed. In other words, a bridging cosolvent $\Phi_{\text{b}}^{\circ}$ contacts two monomers separated by at least three bonds; a mediating cosolvent $\Phi_{\text{m}}^{\circ}$ contacts one monomer and one solvent particle without bridging; and the remainder are passive $\Phi_{\text{p}}^{\circ}$. After classifying the various cosolvents, we report the normalized mean energy per particle for bridging and mediating cosolvents, denoted $\bar{E}_{b}$ and $\bar{E}_{m}$, respectively; the passive set provides a baseline for comparison.
@@ 3 Results and discussion
Our analysis proceeds in two parts. In Section 3.1, we use mean-field theory to identify the energetic regimes that drive cononsolvency through macroscopic phase separation. These results establish conceptual scaffolding and terminology consistent with prior literature. While similar results have appeared variously,34, 50, 64, 51, 67, 52we include this analysis for completeness since application of FHP is new, and specific results here directly inform our subsequent system selection. Readers familiar with cononsolvency may skim this section. In Section 3.2, we employ lattice Monte Carlo simulations to examine microscopic behavior. We strategically select two pairs of systems with equivalent $\chi^{\text{eff}}_{ij}$ from phase-separating regimes, but for a given pair, system differ in their Hamiltonians; one features only isotropic interactions and the other incorporates orientation-dependent terms. Through detailed analysis of coil-globule-coil transitions, which are observed for all systems, we elucidate how interaction anisotropy influences cononsolvency at the microscopic level.
Fig. 2: Classification of phase behavior for ternary polymer solutions described by the FHP framework. (a) Regions in interaction parameter space $\boldsymbol{X}^{\text{FHP}}=(\chi^{\text{FHP}}_{\text{ps}},\chi^{\text{FHP}}_{\text{pc}},\chi^{\text{FHP}}_{\text{sc}})$ where phase separation is required at some composition $\boldsymbol{\phi}$, as determined by the mean-field stability criterion ($|\mathbf{H}|<0$). Markers indicate parameter sets that yield instability; cross-sectional planes at $\chi^{\text{FHP}}_{\text{sc}}=0$, 4-4, 7-7, and 10-10 correspond to panels (b-e). (b-e) Two-dimensional slices of (a) showing phase behavior as a function of $\chi^{\text{FHP}}_{\text{ps}}$ and $\chi^{\text{FHP}}_{\text{pc}}$ for fixed values of $\chi^{\text{FHP}}_{\text{sc}}$. Shaded regions indicate parameter combinations that guarantee phase separation at some composition. Regions of observed two-phase coexistence are labeled $R_{\text{sc}}$, $R_{\text{ps}}$, and $R_{\text{pc}}$, according to the most negative interaction parameter. Square and diamond markers in (b) and (e) denote parameter sets selected for further analysis in Section 3.2.
@@ 3.1 Analysis of phase behavior
Consistent with prior literature using FH theory,34, 50, 64, 51, 67 there are three distinct parameter regimes characterized by the strongest interaction type present (Fig. 2). We denote these regimes as $R_{ij}$, corresponding to the region of parameter space where $(\chi^{\text{FHP}}_{ij}<\chi^{\text{FHP}}_{ik},\chi^{\text{FHP}}_{jk}\leq 0)$ (i.e., the strongest attractive interaction is between species $i$ and $j$). The emergence of cononsolvency, in terms of competing interactions, can be understood by examining successive cross sections in Fig. 2a. Initially, at $\chi^{\text{FHP}}_{\text{sc}}=0$, two distinct regions, $R_{\text{ps}}$ and $R_{\text{pc}}$, exhibit phase separation for a broad range of parameters, for which polymer-solvent or polymer-cosolvent interactions are strongly favorable relative to interactions between solvent and cosolvent (Fig. 2b). As solvent-cosolvent interactions become increasingly favorable (more negative $\chi^{\text{FHP}}_{\text{sc}}$), the areas of $R_{\text{ps}}$ and $R_{\text{pc}}$ shrink, as interactions are more balanced. Concurrently, a third region, $R_{\text{sc}}$, emerges where solvent-cosolvent interactions dominate. (Fig. 2c) Eventually, as solvent-cosolvent affinity strengthens further, $R_{\text{ps}}$ and $R_{\text{pc}}$ vanish entirely, leaving an expanded $R_{\text{sc}}$ region (Fig. 2d,e). This analysis provides a baseline expectation that cononsolvency can be driven by different dominant interactions, be it polymer-solvent, polymer-cosolvent, or solvent-cosolvent, irrespective of microscopic details.
@@
Prior work has suggested that the shape, location, and extent of coexistence regions, as well as the orientation of tie lines, may reflect distinct underlying mechanisms of cononsolvency. 51 To investigate this, we constructed full ternary phase diagrams for representative parameter sets drawn from each of the three regimes: $R_{\text{ps}}$, $R_{\text{pc}}$, and $R_{\text{sc}}$. Indeed, the shapes of binodal curves and the orientations of tie lines differ based on the dominant interaction present in the system (Supplementary Information, Fig. S1) When interactions between a polymer and one of the solvents dominate, tie lines connect phases in which one is rich in polymer and the better solvent, while the other is rich in the lesser solvent and depleted in polymer. When solvent-cosolvent interactions dominate, the tie lines orient nearly perpendicular to the polymer composition axis, leading to coexistence between a very polymer-lean phase with mixed solvent and a polymer-rich phase.
Visually, the orientation of the tielines and the resultant coexistent composition suggest that there is a polymer-lean and polymer-rich phase. To verify that the reported instabilities correspond to cononsolvency, we performed an eigenmode analysis of the Hessian of the free energy at representative compositions within the binodal region. In the $\chi_{\text{pc}}$-dominated regime, the unstable mode corresponds to a channel where the polymer and cosolvent fluctuate together while the solvent fluctuates in the opposite direction. Conversely, in the $\chi_{\text{sc}}$-dominated regime, the instability primarily follows the polymer concentration while solvent and cosolvent fluctuate in phase. These results confirm that cononsolvency arises via distinct mechanisms in the two regimes. Full details of the eigenmode decomposition and channel classification are provided in the Supporting Information (Section S2).
Nevertheless, while these observations may serve as useful indicators of the dominant underlying interactions, they remain qualitative and do not offer insight into the microscopic details of the interactions that underlie cononsolvency.
@@ 3.2 Analysis of simulations
To address this, we perform Monte Carlo (MC) simulations to study the conformational behavior and solvation-shell characteristics of a single polymer chain in mixed solvents for four systems. We present results for a chain with degree of polymerization $N_{\text{m}}=72$; corresponding data for a shorter chain with $N_{\text{m}}=32$ are provided in the SI (Figs. S9–S12). Both lengths show similar physics but the longer chain shows more pronounced effects. In Section 3.2.1, we describe the selection of these systems and broadly categorize their behavior. The results supporting this categorization are then detailed in Sections 3.2.2 and 3.2.3.
@@ 3.2.1 Overview of systems and mechanisms
Table 1: Parameters used in Monte Carlo simulations of FHP models. Parameters for label type $R_{ij}^{\circ}$ are for isotropic simulations $(\Delta_{ij}=0)$ and those for type $R_{ij}^{\curlywedge}$ are for anisotropic simulations $(\Delta_{ij}\neq 0)$.
@@
$R_{\text{sc}}^{\circ}$ -1 -2 -10 0 0
$R_{\text{sc}}^{\curlywedge}$ -1 -2 -10 -0.5 -0.5
$R_{\text{pc}}^{\circ}$ -1 -10 0 0 0
$R_{\text{pc}}^{\curlywedge}$ -1 -10 0 -0.5 -0.5
In all systems, we find evidence of single-chain collapse at certain mixture compositions based on observed $R_{\text{g}}$ relative to that in either pure solvent. Based on further analysis of the solvation environment around the polymer, we then associate these chain collapses with specific mechanisms within the vernacular of the cononsolvency literature. Simulations with parameter sets that are derived from $R_{\text{sc}}$ show a collapse through preferential mixing of solvent and cosolvent; this mechanism tends to yield a “dry” globule. In this case, including anisotropic interactions with cosolvent species qualitatively alters the cosolvent-induced coil-globule transition. Parameter sets from $R_{\text{pc}}$ collapse through preferential adsorption of cosolvent; this mechanism tends to yield a “wet” globule. For this situation, the role of anisotropy is apparent only in nuanced changes to the polymer solvation environment in its collapsed state.
Fig. 3: A comparison of conformational characteristics and environment of a polymer from $R_{\text{sc}}^{\circ}$ (blue circles) and $R_{\text{sc}}^{\curlywedge}$ (orange triangles). (a) Normalized single-chain $R_{\text{g}}$ and normalized excess (b) monomer-monomer, (c) monomer-cosolvent and (d) solvent-cosolvent interactions as cosolvent fraction $x_{c}$ is varied from 0 to 11. In (a), the dashed horizontal line is a guide for the reference to a maximally compact polymer and the white arrow shows the shift in the minima. In (a) and (b), the white arrow shows a marked shift in the behavior of the curves. Horizontal axis labels are share betwen panels (a) and (c) as well as (b) and (d). Error bars correspond to the standard error of the mean and are generally smaller than the symbol size.